# TEn-CATS: Text-Enriched Audio-Visual Video Parsing with Multi-Scale Category-Aware Temporal Graph

### v1

@@ III-D1 Category-Guided Temporal Graph Construction
We aim to construct a category-aware multi-scale temporal graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$ that dynamically models event-specific temporal structures through a learnable category-guided connection. Here, $\mathcal{N}$ is the set of nodes and $\mathcal{E}$ is the set of temporal edges.
In our graph, we set each segment as a node, and the temporal edges between the nodes are dynamically determined based on the predicted event categories. Let $\mathbf{P}\in\mathbb{R}^{b\times T\times C}$ denote the predicted category probabilities for each segment. For each node, we project $\mathbf{P}$ into a set of preference scores $\mathbf{h}\in\mathbb{R}^{b\times T\times K}$ over a candidate hop set $\mathcal{K}=\left\{1,2,...,K\right\}$, via:
@@
where $\mathbf{W}_{K}\in\mathbb{R}^{C\times K}$ is a learnable matrix that maps category probabilities to temporal hop preferences. Then we apply the Gumbel-Softmax [45] sampling to compute the soft selection scores for candidate hops. Gumbel-Softmax is a differentiable approximation to categorical sampling, enabling discrete hop selection while allowing end-to-end gradient-based optimization. Using that, we make the model to dynamically and flexibly select temporal neighbors during training without breaking the differentiability, thus capturing diverse temporal structures more effectively. We compute the hop sampling distribution $s_{t}$ for each:
$s_{t,i}=\frac{exp((h_{t,i}+g_{t,i}))/\tau}{{\textstyle\sum_{j=1}^{K}}exp((h_{t,j}+g_{t,j}))/\tau},s_{t}\in\mathbb{R}^{K}$ (6)
Here $g$ is the the Gumbel noise and $\tau$ is the temperature parameter controlling the sampling sharpness. Based on these scores, we denote the selected hop indices as $\mathcal{L}=\left\{l_{1},l_{2},...,l_{k}\right\}$, Here, $k$ is a hyper-parameter that specifies the number of hops selected from the candidate hop set $\mathcal{K}$.
To further model the temporal sensitivity across different categories, we introduce a learnable decay $\lambda$ computed as:
$\mathbf{\lambda}=\mathbf{PW}_{\lambda}\in\mathbb{R}^{b\times T}$ (7)
$\mathbf{W}_{\lambda}\in\mathbb{R}^{C}$ is a learnable decay vector. The design of $\lambda$ follows the intuition that, in most audio-visual events, the closer the segments are in time, the more likely they are semantically correlated. While the relevance tends to decay as the temporal distance increases. In addition, different categories of events exhibit various temporal sensitivities. For example, transient events such as ”dog barking” may need stronger local correlations, whereas persistent events like ”lawn mower” may require longer temporal dependencies. By learning different decay factors, the model can better adapt to the diverse temporal dynamics of different events.
After selecting the temporal neighbors, we compute the edge weight for each connection. Given a selected hop distance $l_{i}\in\mathcal{L}$ with a soft selection score $s$, we compute the final edge weight using an exponential decay term:
$w_{t,i}=s_{t,i}\times\text{exp}(-{\lambda}_{t}l_{i})$ (8)
This allows the model to assign lower weights to longer-range temporal connections, while retaining differentiability.
In summary, each node not only has self-loop to reserve information of the node, but selects future neighbors based on the sampled hops and decay-modulated weights. The whole process of graph construction we summarized in the Algorithm 1.
@@ III-D2 Multi-Scale Graph Propagation
Following the graph construction described in Section 1), we adopt a multi-scale temporal aggregation network to effectively propagate and refine temporal features over the constructed graph. This component aims to capture hierarchical contextual information while preserving local semantics for each node. As shown in Fig. 2 right, Firstly, we input the node features $\mathbf{X}\in\mathbb{R}^{b\times T\times d}$, and the weighted adjacency matrix $\mathbf{A}$ with edge weight matrix $\mathbf{W}$. Then we perform two layers of residual graph attention to refine the node representations:
$\mathbf{H}^{(1)}=\text{GAT}_{1}(\mathbf{X},\mathbf{A},\mathbf{W})+R_{1}(\mathbf{X})\in\mathbb{R}^{b\times T\times d}$ (9)
$\mathbf{H}^{(2)}=\text{GAT}_{2}(\mathbf{H}^{(1)},\mathbf{A},\mathbf{W})+R_{2}(\mathbf{H}^{(1)})\in\mathbb{R}^{b\times T\times d}$ (10)
$\text{GAT}_{1},\text{GAT}_{2}$ are TransformerConv [46, 47] with 4 attention heads, in which edge weights are passed as one dimensional edge attributes. Specifically, $\mathbf{H}^{(1)}$ and $\mathbf{H}^{(2)}$ are the updated node features after each TransformerConv layer, and $R_{1},R_{2}$ denote the residual mappings to align feature dimensions if necessary.
@@
$\mathbf{X}_{u}=\sigma\!\left(\mathbf{W}_{u}\,[\mathbf{H}^{(2)}\|\mathbf{u}]\right)\odot\mathbf{H}^{(2)}+\Big{(}1-\sigma\!\left(\mathbf{W}_{u}\,[\mathbf{H}^{(2)}\|\mathbf{u}]\right)\Big{)}\odot\mathbf{u}$ (11)
where $\sigma(\cdot)$ denotes the element-wise sigmoid function, which dynamically generates a soft gate between the local and global information for each node. $\mathbf{W}_{u}\in\mathbb{R}^{2d\times d}$ is a learnable linear transformation that projects the concatenated features back to $\mathbb{R}^{b\times T\times d}$. And the operator $\left|\right|$ is the concatenation of features along the channel dimension. In addition, the fused feature $\mathbf{X}_{u}$ is computed as a weighted combination of $\mathbf{H}^{(2)}$ and $\mathbf{u}$, enabling adaptive emphasis on local or global semantics.
Finally, we apply a fully connected layer to obtain the updated node features. Noted that we place the CATS module into audio and visual branch separately and share the parameters with each other.
@@
6:  $\mathcal{E}\leftarrow\mathcal{E}\cup\{(t,t)\}$;  $\mathbf{W}_{t,t}\leftarrow 1$
7:  $s_{t,i}\leftarrow\dfrac{\exp((h_{t,i}+g_{t,i})/\tau)}{\sum_{j=1}^{K}\exp((h_{t,j}+g_{t,j})/\tau)}$, $\forall i\in\mathcal{K}$
@@
9:  $\lambda_{t}\leftarrow\mathbf{P}_{t}\mathbf{W}_{\lambda}$
@@
12:   $\mathbf{W}_{t,\,t{+}l_{i}}\leftarrow s_{t,i}\exp(-\lambda_{t}l_{i})$
13:return $\mathcal{G}=(\mathcal{N},\mathcal{E},\mathbf{W})$
@@ III-E Gated Fusion
To fully exploit the complementary advantages of the feature aggregation and graph reasoning branches, we introduce a gated fusion module that dynamically balances the contributions of HAN-aggregated and CATS-refined features.
We consider that HAN aggregates audio or visual features into temporally-aware semantic representations by emphasizing salient segments via attention mechanisms. This results in $\mathbf{F}_{HAN}\in\mathbb{R}^{b\times T\times d}$, which captures modality-specific semantic importance and local temporal context. Building upon this, CATS module refine the feature representation into $\mathbf{F}_{CATS}\in\mathbb{R}^{b\times T\times d}$ to model long-range dependencies, time-skipped correlations, and event-level structure. Although $\mathbf{F}_{CATS}$ is derived from $\mathbf{F}_{HAN}$, the graph propagation process transforms the feature space by emphasizing structural consistency and relational context. Hence, the two outputs are semantically related but express different views of the input, which are variably useful across segments.
Therefore, we design a gated fusion mechanism to adaptively combine the HAN and CATS outputs at the segment level. Specifically, we first concatenate the features $\mathbf{F}_{HAN}$ and $\mathbf{F}_{CATS}$ along the feature dimension, and use a gating function to calculate the segment-wise fusion weights $\mathbf{G}$:
$\mathbf{G}=\sigma(\mathbf{W}_{f}(\mathbf{F}_{HAN}\left|\right|\mathbf{F}_{CATS})+\mathbf{b}_{f}),\mathbf{G}\in\mathbb{R}^{b\times T\times d}$ (12)
where $\mathbf{W}_{f}\in\mathbb{R}^{2d\times d}$ and $\mathbf{b}_{f}\in\mathbb{R}^{d}$ are the learnable weight matrix and bias of a linear transformation. The final fused representation $\mathbf{F}_{Fuse}$ is computed as:
$\mathbf{F}_{Fuse}=\mathbf{G}\odot\mathbf{F}_{HAN}+(1-\mathbf{G})\odot\mathbf{F}_{CATS}$ (13)
where $\odot$ denotes element-wise multiplication. This gating mechanism enables the model to dynamically favor either the semantic-aware or structure-aware representation at each temporal location, depending on their relative informativeness. It serves as a critical bridge between semantic reasoning and temporal-structural inference in our framework.
@@ IV-A2 Implementation Details
For LLP dataset, we follow VALOR [11] by using the pre-trained CLAP [21] model to extract 768-dimensional audio features from the audio signal. For the visual modality, we employ pre-trained CLIP [22] and 3D ResNet to extract 768-dimensional and 512-dimensional features. In addition, we train our framework using the batch size of 128 for 10 epochs. For the parameters of the CATS module, we use the candidate skip lengths $k\in\{1,2,...,9\}$, where top-3 skip lengths are dynamically selected via Gumbel-Softmax with a temperature of 1.0.
For the UnAV-100 dataset, we follow [48] to extract 2048 dimensional visual features using a two-stream I3D model (RGB + RAFT), and obtain 128 dimensional audio features from pre-trained VGGish aligned with 0.32s video segments [49, 50]. Since UnAV-100 lacks CLIP/CLAP features, we focus solely on evaluating the CATS module. As a self-contained design that does not rely on pseudo labels, this setup enables a fair assessment of CATS’s temporal modeling capacity under large-scale weak supervision. Evaluation of the BiT module is left to the LLP dataset, where high-quality pseudo labels are available.
@@ IV-A4 Baseline Methods
We compare our TEn-CATS with many other baselines to show the effectiveness of our model in two different datasets. For example, we adopt advanced structural methods such as MA [8], CMPAE [18], and CoLeaF [19], and use pseudo-label-based methods such as VALOR [11], LSLD [51], and NREP [13] for comparison.
TABLE I: The performance of TEn-CATS and comparative methods in AVVP, with the best results highlighted in  bold and the second results highlighted in text.
@@ IV-B1 Comparative Performance Overview
As shown in Table I, TEn-CATS achieves significant improvements over prior methods across both segment-level and event-level metrics. Particularly at the segment level, our model demonstrates its superior ability to model fine-grained temporal and semantic structures under weak supervision.
The notable gain in segment-level Event@AV reflects the synergy between our BiT and CATS modules. This shows that the model can propagate consistent semantic signals across time and suppress spurious activations, which is particularly effective when events are temporally fragmented or weak in one modality. Together, these two modules enable the model to identify and localize audio-visual events more accurately at the segment level.
@@ IV-B2 Event-Wise Discriminative Power Performance
To further understand the performance gains of our model, we conduct a per-class event-level F1 score comparison against the baseline CoLeaF [19]. As illustrated in Fig. 4, our model consistently achieves higher per-class F1 scores across nearly all event categories. Notably, substantial improvements are observed in previously underperforming classes in CoLeaF, such as Cheering (from 0.66 to 0.92), Acoustic_guitar (from 0.70 to 0.97), and Telephone_bell_ringing (from 0.80 to 0.92), which are typically characterized by transient and overlapping audio-visual events. The improved recognition of these fine-grained events demonstrates the model’s enhanced ability to disentangle semantically entangled cues.
Despite the fact that most F1 scores of our model exceed 0.90, the metric of event-level Event@AV in Table I is 61.9%. This gap is because Fig. 4 reflects the macro F1, treating each category equally regardless of its frequency. While the Event@AV is micro F1, which is dominated by high-frequency events, such as Speech and Singing. Even slight underperformance in these dominant classes can significantly affect the global metric.

### v2

@@ III-D1 Category-Guided Temporal Graph Construction
We aim to construct a category-aware multi-scale temporal graph $\mathcal{G}=(\mathcal{N},\mathcal{E})$ that dynamically models event-specific temporal structures through a learnable category-guided connection, where $\mathcal{N}$ is the set of nodes and $\mathcal{E}$ is the set of temporal edges. In this graph, we set each segment as a node, and the temporal edges between the nodes are dynamically determined based on the predicted event categories. Let $\mathbf{P}\in\mathbb{R}^{B\times T\times C}$ denote the predicted category probabilities for each segment. For each node, we project $\mathbf{P}$ into a set of preference scores $\mathbf{H}\in\mathbb{R}^{B\times T\times K}$ over a candidate hop set $\mathcal{K}=\left\{1,2,...,K\right\}$, via:
@@
where $\mathbf{W}_{K}\in\mathbb{R}^{C\times K}$ is a learnable matrix that maps category probabilities to temporal hop preferences. Then we apply the Gumbel-Softmax [42] sampling to compute the soft selection scores for candidate hops. Gumbel-Softmax is a differentiable approximation to categorical sampling, enabling discrete hop selection while allowing end-to-end gradient-based optimization. Using that, the model is enabled to dynamically and flexibly select temporal neighbors during training without breaking the differentiability, thus capturing diverse temporal structures more effectively. We compute the hop sampling distribution $s_{t}$ for each:
$\mathbf{s}_{t,i}=\frac{exp((\mathbf{h}_{t,i}+\mathbf{n}_{t,i}))/\tau}{{\textstyle\sum_{j=1}^{K}}exp((\mathbf{h}_{t,j}+\mathbf{n}_{t,j}))/\tau}\in\mathbb{R}^{B}$ (6)
where $\mathbf{n}_{t,i}$ is the Gumbel noise and $\tau$ is the temperature parameter controlling the sampling sharpness. Based on these scores, we denote the selected hop indices as $\mathcal{L}=\left\{l_{1},l_{2},...,l_{k}\right\}$, Here, $k$ is a hyper-parameter that specifies the number of hops selected from the candidate hop set $\mathcal{K}$.
To further model the temporal sensitivity across different categories, we introduce a learnable decay $\mathbf{Q}$ computed as:
$\mathbf{Q}=\mathbf{Pw}_{\lambda}\in\mathbb{R}^{B\times T}$ (7)
where $\mathbf{w}_{\lambda}\in\mathbb{R}^{C}$ is a learnable decay vector. The design of $\mathbf{Q}$ follows the intuition that, in most audio-visual events, the closer the segments are in time, the more likely they are semantically correlated, although such a correlation tends to decay as the temporal distance increases. In addition, different events exhibit different temporal sensitivities. For example, “dog barking” is often a transient event, with strong local correlations, while “lawn mower” is a persistent event, with long temporal dependencies. By learning different decay factors, the model can better adapt to the temporal dynamics of different events.
After selecting the temporal neighbors, we compute the edge weight for each connection. Given a selected hop distance $l_{i}\in\mathcal{L}$ with a soft selection score $\mathbf{s}$, we compute the final edge weight using an exponential decay term:
$\mathbf{w}_{t,i}=\mathbf{s}_{t,i}\odot\text{exp}(-{\mathbf{q}}_{t}l_{i})$ (8)
where $\odot$ denotes element-wise multiplication and $\mathbf{q}_{t}\in\mathbb{R}^{B}$ is the $t$-th column vector of $\mathbf{Q}$. This allows the model to assign lower weights to longer-range temporal connections, while retaining differentiability.
In summary, each node not only has self-loop to reserve information of the node, but selects future neighbors based on the sampled hops and decay-modulated weights. The whole process of graph construction is summarized in Algorithm 1.
@@ III-D2 Multi-Scale Graph Propagation
Following the graph construction described in Section III-D1, we adopt a multi-scale temporal aggregation network to effectively propagate and refine temporal features over the constructed graph. This component aims to capture hierarchical contextual information while preserving local semantics for each node. As shown in Fig. 2, firstly, we input the node features $\mathbf{X}\in\mathbb{R}^{B\times T\times d}$, and the weighted adjacency matrix $\mathbf{A}$ with edge weight matrix $\mathbf{W}$ with elements $\mathbf{w}_{t,i}$ calculated in 8. Then, we perform two layers of residual graph attention to refine the node representations:
$\mathbf{Z}^{(1)}=\text{GAT}_{1}(\mathbf{X},\mathbf{A},\mathbf{W})+R_{1}(\mathbf{X})\in\mathbb{R}^{b\times T\times d}$ (9)
$\mathbf{Z}^{(2)}=\text{GAT}_{2}(\mathbf{Z}^{(1)},\mathbf{A},\mathbf{W})+R_{2}(\mathbf{Z}^{(1)})\in\mathbb{R}^{b\times T\times d}$ (10)
where $\text{GAT}_{1}$ and $\text{GAT}_{2}$ are TransformerConv [43, 44] with four attention heads, in which the edge weights $\mathbf{W}\in\mathbb{R}^{b\times T\times T}$ are passed as one dimensional edge attributes. Specifically, $\mathbf{Z}^{(1)}$ and $\mathbf{Z}^{(2)}$ are the updated node features after each TransformerConv layer, and $R_{1}$ and $R_{2}$ denote the residual mappings to align the feature dimensions if necessary.
@@
$\mathbf{\tilde{X}}=\sigma\!\left(\mathbf{\tilde{W}}\,[\mathbf{Z}^{(2)}\|\mathbf{U}]\right)\odot\mathbf{Z}^{(2)}+\Big(1-\sigma\!\left(\mathbf{\tilde{W}}\,[\mathbf{Z}^{(2)}\|\mathbf{U}]\right)\Big)\odot\mathbf{U}$ (11)
where $\sigma(\cdot)$ denotes the element-wise sigmoid function, which dynamically generates a soft gate between the local and global information for each node. $\mathbf{\tilde{XW}}\in\mathbb{R}^{2d\times d}$ is a learnable linear transformation that projects the concatenated features back to $\mathbb{R}^{b\times T\times d}$. The operator $\left|\right|$ is the concatenation of features along the channel dimension. In addition, the fused feature $\mathbf{\tilde{X}}$ is computed as a weighted combination of $\mathbf{H}^{(2)}$ and $\mathbf{U}$, enabling adaptive emphasis on local and global semantics.
Finally, we apply a fully connected layer to obtain the updated node features. Noted that we place the CATG module into audio and visual branch separately and share the parameters with each other.
@@
6:  $\mathcal{E}\leftarrow\mathcal{E}\cup\{(t,t)\}$;  $\mathbf{w}_{t,t}\leftarrow\mathbf{1}\in\mathbb{R}^{B}$}
7:  $\mathbf{s}_{t,i}\leftarrow\dfrac{\exp((\mathbf{h}_{t,i}+\mathbf{n}_{t,i})/\tau)}{\sum_{j=1}^{K}\exp((\mathbf{h}_{t,j}+\mathbf{n}_{t,j})/\tau)}$, $\forall i\in\mathcal{K}$
@@
9:  $\mathbf{q}_{t}\leftarrow\mathbf{P}_{t}\mathbf{w}_{\lambda}$
@@
12:   $\mathbf{w}_{t,\,i}\leftarrow\mathbf{s}_{t,i}\odot\exp(-\mathbf{q}_{t}l_{i})$
13:Assemble the adjacency matrix $\mathbf{W}$ where each entry $\mathbf{W}_{b,\,t,\,t{+}l_{i}}=\mathbf{w}_{t,i}$, and $\mathbf{w}_{t,i}$ is computed using the learnable parameters $\mathbf{W}_{K}$ and $\mathbf{w}_{\lambda}$ as:
$\mathbf{w}_{t,i}=\mathrm{Softmax}\!\left(\mathbf{P}_{t}\mathbf{W}_{K}\right)\cdot\exp(-\mathbf{w}_{\lambda}\,l_{i}),\quad\mathbf{W}\in\mathbb{R}^{B\times T\times T}.$
14:Optimization: Gradients of $\mathbf{W}_{K}$ and $\mathbf{w}_{\lambda}$ are obtained through back-propagation from the overall task loss $\mathcal{L}_{\text{task}}$ of our network (since the graph construction is fully differentiable in Eqs. (5)–(8)). Both parameters are updated via gradient descent:
$\mathbf{W}_{K}\leftarrow\mathbf{W}_{K}-\eta\frac{\partial\mathcal{L}_{\text{task}}}{\partial\mathbf{W}_{K}},\quad\mathbf{w}_{\lambda}\leftarrow\mathbf{w}_{\lambda}-\eta\frac{\partial\mathcal{L}_{\text{task}}}{\partial\mathbf{w}_{\lambda}}.$
15:return $\mathcal{G}=(\mathcal{N},\mathcal{E},\mathbf{W})$
@@ III-E Gated Fusion
To fully exploit the complementary advantages of the feature aggregation and graph reasoning branches, we introduce a gated fusion module that dynamically balances the contributions of HAN-aggregated and CATG-refined features.
We consider that HAN aggregates audio or visual features into temporally-aware semantic representations by emphasizing salient segments via attention mechanisms. This results in $\mathbf{M}_{HAN}\in\mathbb{R}^{b\times T\times d}$, which captures modality-specific semantic importance and local temporal context. Building upon this, CATG module refine the feature representation into $\mathbf{M}_{CATG}\in\mathbb{R}^{b\times T\times d}$ to model long-range dependencies, time-skipped correlations, and event-level structure. Although $\mathbf{M}_{CATG}$ is derived from $\mathbf{M}_{HAN}$, the graph propagation process transforms the feature space by emphasizing structural consistency and relational context. Hence, the two outputs are semantically related but express different views of the input, which are variably useful across segments.
Therefore, we design a gated fusion mechanism to adaptively combine the HAN and CATG outputs at the segment level. Specifically, we first concatenate the features $\mathbf{M}_{HAN}$ and $\mathbf{M}_{CATG}$ along the feature dimension, and use a gating function to calculate the segment-wise fusion weights $\mathbf{G}$:
$\mathbf{O}=\sigma\big(\mathbf{\bar{W}}[\mathbf{M}_{HAN};\mathbf{M}_{CATG}]+\mathbf{\bar{b}}\big),\quad\mathbf{O}\in\mathbb{R}^{b\times T\times d}$ (12)
where $\mathbf{\bar{W}}\in\mathbb{R}^{2d\times d}$ and $\mathbf{\bar{b}}\in\mathbb{R}^{d}$ are the learnable weight matrix and bias of a linear transformation. The final fused representation $\mathbf{F}_{Fuse}$ is computed as:
$\mathbf{M}_{Fuse}=\mathbf{O}\odot\mathbf{M}_{HAN}+(\mathbf{1}-\mathbf{O})\odot\mathbf{M}_{CATG}$ (13)
where $\odot$ denotes element-wise multiplication, and $\mathbf{1}$ is a tensor of the same dimension as $\mathbf{O}$, with all elements equal 1. This gating mechanism enables the model to dynamically favor either the semantic-aware or structure-aware representation at each temporal location, depending on their relative informativeness. It serves as a critical bridge between semantic reasoning and temporal-structural inference in our framework.
@@ IV-A2 Implementation Details
For the LLP dataset, we follow VALOR [9] by using the pre-trained CLAP [14] model to extract 768-dimensional audio features from the audio signal. For the visual modality, we employ pre-trained CLIP [15] and 3D ResNet to extract 768-dimensional and 512-dimensional features. In addition, we train our framework using the batch size of 128 for 10 epochs. For the parameters of the CATG module, we use the candidate skip lengths $k\in\{1,2,...,9\}$, where top-3 skip lengths are dynamically selected via Gumbel-Softmax with a temperature of 1.0.
For the UnAV-100 dataset, we follow [45] to extract 2048 dimensional visual features using a two-stream I3D model (RGB + RAFT), and obtain 128 dimensional audio features from pre-trained VGGish aligned with 0.32s video segments [46, 47]. Since UnAV-100 lacks CLIP/CLAP features, we focus solely on evaluating the CATG module. As a self-contained design that does not rely on pseudo labels, this setup enables a fair assessment of CATG’s temporal modeling capacity under large-scale weak supervision. Evaluation of the BiT module is left to the LLP dataset, where high-quality pseudo labels are available.
@@ IV-A4 Baseline Methods
To show the effectiveness of our model, we compare our TEn-CATG with several other baselines on two different datasets, including advanced structural methods such as MA [6], CMPAE [18], and CoLeaF [19], and pseudo-label-based methods such as VALOR [9], LSLD [48], and NREP [11].
TABLE I: The performance of TEn-CATG and comparative methods in AVVP, with the best results highlighted in bold and the second results highlighted in text.
@@ IV-B1 Comparative Performance Overview
As shown in Table I, TEn-CATG achieves significant improvements over prior methods across both segment-level and event-level metrics. Particularly at the segment level, our model demonstrates its superior ability to model fine-grained temporal and semantic structures under weak supervision.
The notable gain in segment-level Event@AV reflects the synergy between our BiT and CATG modules. This shows that the model can propagate consistent semantic signals across time and suppress spurious activations, which is particularly effective when events are temporally fragmented or weak in one modality. Together, these two modules enable the model to identify and localize audio-visual events more accurately at the segment level.
@@ IV-B2 Event-Wise Discriminative Power Performance
To further understand the performance gains of our model, we conduct a per-class event-level F1 score comparison against the baseline CoLeaF [19]. As illustrated in Fig. 4, our model consistently achieves higher per-class F1 scores across nearly all event categories. Notably, substantial improvements are observed in previously underperforming classes in CoLeaF, such as Cheering (from 0.66 to 0.92), Acoustic_guitar (from 0.70 to 0.97), and Telephone_bell_ringing (from 0.80 to 0.92), which are typically characterized by transient and overlapping audio-visual events. The improved recognition of these fine-grained events demonstrates the model’s enhanced ability to disentangle semantically coupled cues.
Despite the fact that most F1 scores achieved by our model exceed 0.90, the metric of event-level Event@AV in Table I is 61.9%. This gap is because Fig. 4 reflects the macro F1, treating each category equally regardless of its frequency. While the Event@AV is micro F1, which is dominated by high-frequency events, such as Speech and Singing. Even slight underperformance in these dominant classes can significantly affect the global metric.